How much of a 25% alloy and how much of a 50% alloy should be mixed to make 1000g of a 45% alloy
I have tried:
.25x+.5y=.45(1000)
x+y=1000
y=3000 and X=-2000
Answer doesn't make sense

Found 2 solutions by nerdybill, scott8148:
Answer by nerdybill(7384)

How much of a 25% alloy and how much of a 50% alloy should be mixed to make 1000g of a 45% alloy
Let x = amount (g) of 25% alloy
then
1000-x = amount(g) of 50% alloy
.
.25x + .50(1000-x) = .45(1000)
.25x + 500-.50x = 450
-.25x + 500 = 450
-.25x = -50
x = 200 g (25% alloy)
.
50% alloy:
1000-x = 1000-200 = 800 g

Answer by scott8148(6628)

your equations are fine

multiplying 1st eqn by -4 ___ -x - 2y = -1800

adding 2nd eqn to eliminate x ___ -x - 2y + x + y = -1800 + 1000 ___ -y = - 800

solve for y; then substitute back to find x